# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

### v2

@@ IV-C Experiments
This section explains how we conduct our experiments to evaluate TEASMA and answer the research questions described in Section IV-A. To answer RQ1, we perform the first step of the TEASMA approach (Section III) as detailed in Algorithm 1 using the training input set of our subjects. As a result, we calculate FDR and MS for a large number of diverse input subsets sampled from the training set. Subsequently, we measure the correlation between FDR and MS using the non-parametric Spearman coefficient based on all subsets, since we cannot assume linearity. We also measure the correlation for mutants generated by each MO separately and then attempt to find subsets of operators leading to higher correlation. Before proceeding to the next step, we compare the highest correlation achieved by subsets of MOs with that obtained with all MOs. If the correlation of the former is higher than the latter, then we only consider mutants generated by the optimal subset with the highest correlation. Next, using all the training subsets, we proceed to the second step of TEASMA, regression analysis, relying on both linear and non-linear models. TEASMA performs a $K$-fold cross-validation procedure with $K=5$ and reports the average of $R^{2}$, MMRE, and RMSE across all folds for each regression model separately. The best regression model can be selected based on these metrics.
@@
To summarize, the experiment that we conducted using the training set to answer RQ1 included the following steps:
@@
    Perform the first step of TEASMA and calculate MS and FDR for each input subset
@@
    Measure the Spearman correlation between FDR and MS
@@
    Perform the second step of TEASMA and conduct regression analysis between FDR and MS using linear and non-linear models (quadratic, exponential, and regression trees)
@@
To address RQ2, we sample a large number of input subsets from the test set, calculate their MS using previously generated mutants, and leverage the previously built and selected regression model to predict FDR ($\widehat{FDR}(T)$). We should note that we determine killed mutants by comparing their output with that of the original model and thus the actual labels of test inputs are not required. To estimate the actual FDR ($ActualFDR$) of each test subset, we use the same fault clusters identified based on the training set as described in Section III-B and we assign each mispredicted test input to one of the clusters. Since the fault estimation approach (Section II-E) employs the HDBSCAN algorithm [27], a density-based hierarchical clustering method where each cluster is characterized by a number of core points, we assign each mispredicted test input to the cluster featuring a core point that is closest to the input. To ensure an accurate calculation of $ActualFDR$ and allow the test set to reach the maximum FDR value of 1, we exclusively take into account fault clusters that can be detected by the test set. To achieve this, we limit our denominator in Equation (10) to the number of fault clusters with at least one mispredicted test input assigned to them. Consequently, we conduct our experiments based on the faults detectable by the test set, so that our results are not affected by the varying quality of the provided test sets across subjects.
Subsequently, we are able to compare for each test subset, the predicted FDR with the actual one and investigate if the former is closely aligned with the latter and can thus be trusted to make decisions about the test set. Therefore, to answer RQ2, we went through the following steps:
@@
    Calculate MS and FDR ($ActualFDR$) for each test subset
@@
    Use previously selected prediction model to compute $\widehat{FDR}$ based on MS for each test subset
@@
As described in Section II-B, different methods have been proposed to determine whether a mutant is killed and therefore to calculate MS. To fully address our research questions, we ran a comprehensive analysis by considering all the MS calculation variants described in Section II-B. As a result, we ran three variations of our experiments that only differ in the way they calculate MS: (E1) $DeepMutationMS$, (E2) $StandardMS$, and (E3) $KSBasedMS$.
Experiments on all subjects took about two days to run using a cloud computing environment provided by the Digital Research Alliance of Canada [32], on the Cedar cluster with a total of 1352 NVIDIA P100 Pascal GPUs with 8 to 16 GB of memory.
V Data Availability
@@
The replication package for our studies will be shared upon the paper’s acceptance.
@@
TABLE III: Spearman correlation coefficients computed between MS and FDR for random input subsets of the training set for each subject
Experiment Subjects
Variants S1 S2 S3 S4 S5 S6
E1 0.99 0.99 0.99 0.99 0.99 0.98
E2 0.92 0.73 0.51 0.91 0.86 0.91
@@
Figure 3: Selected regression models across subjects based on the training set
VI Results
In this section, we report results pertaining to our research questions and discuss their practical implications.
VI-A RQ1: Can an accurate regression model be built to explain FDR as a function of MS?
To answer this question, we first investigate the extent of the correlation between MS and FDR. We perform our experiment on a large number of randomly and uniformly sampled input subsets from the training set, also considering different MS calculation variants. The correlation analysis results using Spearman correlation coefficients for experiment variants E1 and E2 are shown in Table III. We only report results for random subsets as they yielded slightly higher correlations and reporting results from uniform subsets does not provide new insights. The best correlation achieved for each subject across experiment variants, given in each column of Table III, is highlighted in bold font. We did not report the results for experiment variant E3 since all of the correlation values for both random and uniform subsets across all subjects are below 0.1, indicating an absence of correlation. This can be explained by the fact that MS is determined by computing the average $KillingScore$ (Equation (3)) of all inputs within a subset. Consequently, for subsets of different sizes, MS calculated based on $KSBasedMS$ (Equation (4)) falls within a similar range (0 to 0.1). As discussed in Section II-B, this confirms that calculating the MS of a subset as the average capability of its inputs to kill mutants cannot be used as an indicator of the capability of the entire subset to kill mutants.
E1 shows the strongest correlation across all subjects between MS and FDR (above 0.98 across all subjects). In this experiment, we calculate MS as proposed originally by DeepMutation (Equation (2)). In contrast, for experiment E2, $R^{2}$ ranges from 0.51 to 0.92, which is still a strong correlation. The method for identifying killed mutants and calculating MS therefore has a substantial impact on the results of mutation analysis. Given the strong correlations achieved in E1 and E2 across all subjects, we conclude that mutants generated by DeepMutation’s post-training MOs are highly associated with faults. Given the above results, we focus solely on $DeepMutationMS$ and random subsets from now on. Before proceeding to the next steps, we calculated the correlation between FDR and MS individually for each MO as well as for different subsets of operators. However, none of them achieved a significantly higher correlation than the correlation obtained with all MOs. Consequently, we proceeded to the next step considering, for each subject, all of the applicable MOs and performed regression analysis to model the relationship between MS and FDR on the training set. For each subject, we examined multiple regression models starting with a simple linear regression model and then using more complex non-linear models (quadratic, exponential, and regression trees) when needed. We then performed a 5-fold cross-validation procedure to calculate $R^{2}$, MMRE, and RMSE and selected the most accurate model for each subject. Due to space limitation, we only report the results for the selected regression model for each subject in Table IV. These results show a minimum $R^{2}$ value of 0.98 across all subjects, indicating that MS explains most of the variance in FDR. The average RMSE across all $K$-folds for all subjects ranges from 0.03 to 0.05, highlighting the high prediction accuracy. Additionally, the average MMRE varies between 0.1 and 0.18 indicating a low relative error. This suggests that MS can be used to very accurately predict FDR in the training set.
TABLE IV: Accuracy and shape of the selected regression models when predicting FDR with MS, based on the training set and using a 5-fold cross-validation procedure
Accuracy Subjects
metric S1 S2 S3 S4 S5 S6
$R^{2}$ 0.99 0.99 0.98 0.98 0.99 0.98
MMRE 0.18 0.13 0.15 0.14 0.1 0.15
RMSE 0.03 0.04 0.04 0.05 0.03 0.04
Shape* RT RT RT E RT RT
*Selected regression model shapes include Exponential (E) and Regression Tree (RT)
@@ VI-A RQ1: Can an accurate regression model be built to explain FDR as a function of MS?
The final row of Table IV reports the shape of the selected regression model on each subject. Regression trees are the best for all subjects but S4. This suggests that standard regression functions are often not a good fit and regression trees offer more flexibility by optimally partitioning the MS range and computing averages within partitions. To prevent overfitting of the regression trees, which would yield poor results on test sets, we limited their maximum depth to 5.
@@
To visualize our results, we present for each subject in Figure 3, the scatterplots and fitted regression curves between FDR and MS along with 95% prediction intervals (PI), based on all input subsets sampled from the training set. We compute PIs for regression tree models based on a bootstrapping method for non-parametric regression models [33]. PIs depict, for each specific MS value, the likely range of potential FDR values for a 95% confidence level. The narrow PIs across all subjects suggest that the regression models can be confidently used to predict the FDR of a test set with limited uncertainty. We can also observe that the relationship is clearly not linear in most cases. For all subjects, as subset size increases, MS initially shows rapid growth compared to FDR, indicating that the first mutants are relatively easy to kill when compared to faults. But then the rate of increase in MS starts to quickly slow down over time, suggesting that mutants are rapidly becoming more difficult to kill. This confirms that many mutants are not comparable to faults but that nevertheless does not prevent us from accurately modeling their relationship. The highly non-linear relationship between MS and FDR on most subjects indicates that MS cannot be directly used to assess test sets and decide on their adequacy. As illustrated in Figure 3, identical MS values can result in very different levels of fault detection abilities across subjects. For instance, in subject S6, an MS of 0.8 is associated with an FDR of 0.2 while in subject S4, the same MS is associated with an FDR of 0.8.
@@
It should be noted that the shape of the relationship between MS and FDR greatly varies across subjects, which is likely due to the characteristics of the DNN models and associated training sets. For example, differences in internal architectures across DNN models determine the distribution of mutants generated by post-training MOs and contribute to such variation in relationships. Though this is not an issue in practice, this highlights the need to determine the best regression model empirically for each DNN model based on its training set, as specified by TEASMA (Section III).
@@
Answer to RQ1: Post-training MOs generate mutants that are highly associated with faults; thus, based on these mutants, MS has a very strong correlation with FDR, especially when computed with $DeepMutationMS$. Further, an accurate regression model with a minimum $R^{2}=0.98$ (based on cross-validation) can be fitted for each subject based on its training set to accurately capture the relationship between FDR and MS. The shape of the regression model greatly varies across subjects.
VI-B RQ2: Can a test set’s FDR be accurately predicted from its MS using the best regression model built on the training set?
To address this question, we analyze the relationship of the predicted FDR ($\widehat{FDR}$) with the actual one ($ActualFDR$) in the test set. To do so, for each subject, we selected a large number of random test subsets and measured the fit ($R^{2}$) and slope of a linear regression model between $\widehat{FDR}$ and $ActualFDR$. If the former can be trusted as an indicator of the latter, we should have a high $R^{2}$ and a slope close to 1.
@@
TABLE V: Linear correlation between $\widehat{FDR}$ and $ActualFDR$ of all random test subsets across subjects
Accuracy Subjects
metrics S1 S2 S3 S4 S5 S6
$R^{2}$ 0.97 0.98 0.94 0.97 0.98 0.95
Slope 1.21 1.19 1.16 1.23 0.95 1.15
Table V reports the results, where we can see a minimum $R^{2}$ value of 0.94 and a slope that is either slightly below or above 1, across all subjects. Such results indicate that the regression models developed on the training set yield FDR predictions that are closely aligned with the actual FDR of the test sets. Therefore, FDR predictions provide a good basis on which to assess the adequacy of test sets and whether their results can be trusted to decide about deploying a DNN model.
In Figure 4, we depict the linear regression lines between $\widehat{FDR}$ and $ActualFDR$ where the intercept is set to zero, clearly showing a strong alignment between the two. The slopes of the lines fall within the range [0.95 1.23], thus indicating no strong bias in the FDR predictions . In most subjects, the prediction models tend to slightly underestimate the actual FDR, with the exception of subject S5, where an overestimation is observed.
Answer to RQ2: $\widehat{FDR}$ is strongly correlated and therefore closely aligned with $ActualFDR$, thus indicating that, for each subject, the selected regression model built on the training set accurately predicts the FDR of a test set based on its MS. Consequently, TEASMA provides a reliable basis for assessing test sets and whether to trust their results when validating a DNN model.

### v3

@@ IV-D Experiments
This section explains how we conduct our experiments to evaluate TEASMA individually for each selected metric and answer the research questions described in Section IV-A. To answer RQ1, we perform the first step of the TEASMA approach (Section III) as detailed in Algorithm 1 using the training input set of our subjects. As a result, we calculate FDR and AS for a large number of diverse input subsets sampled from the training set. Subsequently, we measure the correlation between FDR and AS using the non-parametric Spearman coefficient based on all subsets, since we cannot assume linearity. Next, using all the training subsets, we proceed to the second step of TEASMA, regression analysis, relying on both linear and non-linear models. TEASMA performs a $K$-fold cross-validation procedure with $K=5$ and reports the average of $R^{2}$, MMRE, and RMSE across all folds for each regression model separately. The best regression model can be selected based on these metrics.
@@
To summarize, the experiment that we conducted for each adequacy metric using the training set to answer RQ1 included the following steps:
@@
    Perform the first step of TEASMA and calculate AS and FDR for each input subset
@@
    Measure the Spearman correlation between FDR and AS
@@
    Perform the second step of TEASMA and conduct regression analysis between FDR and AS using linear and non-linear models (quadratic, exponential, and regression trees)
@@
To address RQ2, we sample a large number of input subsets from the test set, calculate their AS using the same adequacy metric and configuration from the previous steps, and leverage the previously built and selected regression model to predict FDR ($\widehat{FDR}(T)$). We should note that for calculating MS, we determine killed mutants by comparing their output with that of the original model and thus the actual labels of test inputs are not required. Similarly, calculating AS based on SC and IDC does not necessitate knowledge of the inputs’ labels. It is important to remark that we consistently use the same set of randomly sampled test subsets in our experiments with all adequacy metrics.
To estimate the actual FDR ($ActualFDR$) of each test subset, we use the same fault clusters identified based on the training set as described in Section III-B and we assign each mispredicted test input to one of the clusters. Since the fault estimation approach (Section II-D) employs the HDBSCAN algorithm [30], a density-based hierarchical clustering method where each cluster is characterized by a number of core points, we assign each mispredicted test input to the cluster featuring a core point that is closest to the input. To ensure an accurate calculation of $ActualFDR$ and allow the test set to reach the maximum FDR value of 1, we exclusively take into account fault clusters that can be detected by the test set. To achieve this, we limit our denominator in Equation (6) to the number of fault clusters with at least one mispredicted test input assigned to them. Consequently, we conduct our experiments based on the faults detectable by the test set, so that our results are not affected by the varying quality of the provided test sets across subjects.
Subsequently, we are able to compare for each test subset, the predicted FDR with the actual one and investigate if the former is closely aligned with the latter and can thus be trusted to make decisions about the test set. Therefore, to answer RQ2 for each metric, we went through the following steps:
@@
    Calculate AS, using the same metric and configuration from the previous steps, and FDR ($ActualFDR$) for each test subset
@@
    Use previously selected prediction model to compute $\widehat{FDR}$ based on AS for each test subset
@@
To address RQ3, we investigate the accuracy of selected metrics within TEASMA, in in terms of FDR predictions. Specifically, we focus on adequacy metrics identified in RQ2 that exhibit better FDR predictions characterized by higher linear correlation ($R^{2}$) with the actual FDR and slopes closer to 1. We compute RMSE between $\widehat{FDR}$ and $ActualFDR$ and compare such prediction accuracy for each metric. Subsequently, we identify the metric with the lowest average RMSE across subjects, to determine the best metric.
@@
As described in Section II-A3, different methods have been proposed to determine whether a mutant is killed and therefore to calculate MS. To fully address our research questions, we ran a comprehensive analysis by considering all the MS calculation variants described in Section II-A3. As a result, we ran three variations of our experiments that only differ in the way they calculate MS: (E1) $DeepMutationMS$, (E2) $StandardMS$, and (E3) $KSBasedMS$.
The total execution time to conduct our entire experiments, on all subjects using the four investigated adequacy metrics, on one Nvidia P100 GPU with 12GB memory, would have taken approximately five months. Through extensive parallelization and using two separate computing environments, we reduced it to 40 days. We conducted our experiments with the IDC metric on a machine with two Nvidia Quadro RTX 6000 GPUs (each with 24GB GPU memory), an Intel Xeon Gold 6234 16-Core CPU, and 187GB of RAM. For the remaining metrics, we used a cloud computing environment provided by the Digital Research Alliance of Canada [47], utilizing the Cedar cluster with 1352 various Nvidia GPUs (P100 Pascal and V100 Volta) with memory ranging from 8 to 32 GB.
IV-E Data Availability
@@
The replication package for our studies will be shared upon the paper’s acceptance [44].
V Results
In this section, we report results pertaining to our research questions and discuss their practical implications. Furthermore, to enable a full validity analysis of our results, we report the details of the fault clusters used in our experiments.
V-A RQ1: Can an accurate regression model be built to explain FDR as a function of AS, based on any of the existing adequacy metrics?
V-A1 RQ1.1
Which adequacy metrics have a strong correlation with FDR?
To investigate the extent of the correlation between AS and FDR, for each adequacy metric, we randomly sample a large number of input subsets from the training set. We then compute the non-parametric Spearman correlation coefficient, for all adequacy metrics in Table II, between AS and FDR for all these subsets.
@@ V Results
TABLE II: Spearman correlation coefficients computed between AS and FDR using each adequacy metric, for random input subsets of the training set for each subject
Adequacy metric Subjects
Variants S1 S2 S3 S4 S5 S6 S7 S8
MS E1 0.99 0.99 0.99 0.99 0.99 0.98 1 1
E2 0.92 0.73 0.66 0.91 0.86 0.91 0.53 0.45
SC DSC 0.99 0.99 0.99 0.99 0.99 0.99 0.99 1
LSC 0.99 0.99 0.99 0.99 0.99 0.99 0.99 1
IDC 0.99 0.99 0.99 0.99 0.99 1 1 1
When using MS for calculating AS, we consider different MS calculation variants. The correlation analysis results using Spearman correlation coefficients for experiment variants E1 and E2 are shown in Table II. We further extended our experiment to include a large number of uniformly sampled input subsets from the training set since the MS calculation proposed by DeepMutation [2] in Equation (2) considers the number of classes of each mutant killed by inputs in a test set. However, we only report results for random subsets as they yielded slightly higher correlations and reporting results from uniform subsets does not provide new insights.
We did not report the results for experiment variant E3 since all of the correlation values for both random and uniform subsets across all subjects are below 0.1, indicating an absence of correlation. This can be explained by the fact that MS is determined by computing the average $KillingScore$ (Equation (3)) of all inputs within a subset. Consequently, for subsets of different sizes, MS calculated based on $KSBasedMS$ (Equation (4)) falls within a similar range (0 to 0.1). As discussed in Section II-A3, this confirms that calculating the MS of a subset as the average capability of its inputs to kill mutants cannot be used as an indicator of the capability of the entire subset to kill mutants.
E1 shows the strongest correlation across all subjects between MS and FDR (above 0.98 across all subjects). In this experiment, we calculate MS as proposed originally by DeepMutation (Equation (2)). In contrast, for experiment E2, $R^{2}$ ranges from 0.45 to 0.92, which is still a strong correlation but not nearly as high as for E1. The method for identifying killed mutants and calculating MS therefore has a substantial impact on the results of mutation analysis. Given the strong correlations achieved in E1 and E2 across all subjects, we conclude that mutants generated by DeepMutation’s post-training MOs are highly associated with faults. Given the above results, in our experiments using MS for AS calculation, we focus solely on $DeepMutationMS$ and random subsets from now on.
As reported in Table II, the correlation analysis using Spearman correlation coefficients for DSC, LSC, and IDC showed results similar to MS based on variant E1, indicating a strong correlation (above 0.99) across all subjects. These results suggest that all utilized metrics, DSC, LSC, IDC, and MS calculated using $DeepMutationMS$ are all promising metrics for test assessment. However, beyond demonstrating a strong correlation, it is essential to investigate the shape of the relationship between AS and FDR for these metrics, which we explore in the following section.
Before proceeding to the next steps with MS, we investigated the potential impact of MO choices when calculating MS. Ideally, MOs capable of generating mutants associated with faults are preferred in mutation analysis [17]. Therefore, it is imperative to select mutants based on which MS exhibits a stronger correlation with FDR. We calculated the correlation between FDR and MS individually for each MO as well as for different subsets of operators. Our results show that none of the MOs or MO subsets achieved a significantly higher correlation than the correlation obtained with all MOs. Consequently, we proceeded to the next step considering, for each subject, all of the applicable MOs and performed regression analysis to model the relationship between MS and FDR on the training set. Nevertheless, to provide comprehensive results, we reported the correlation for each individual MO and all MO subsets in our replication package [44].
Answer to RQ1.1: Distribution-based coverage metrics including DSC, LSC, and IDC have a very strong Spearman correlation with FDR. Post-training MOs generate mutants that are highly associated with faults; thus, based on these mutants, MS also has a very strong correlation with FDR, especially when computed with $DeepMutationMS$.
TABLE III: Accuracy and shape of the selected regression models, built using each adequacy metric, when predicting FDR based on the training set and using a 5-fold cross-validation procedure
Adequacy Accuracy Subjects
metric metric S1 S2 S3 S4 S5 S6 S7 S8
MS (E1) $R^{2}$ 0.99 0.99 0.99 0.98 0.99 0.98 0.99 1
MMRE 0.18 0.13 0.16 0.14 0.1 0.13 0.25 0.04
RMSE 0.03 0.04 0.04 0.05 0.03 0.05 0.03 0.01
Shape* RT RT RT E RT RT L RT
DSC $R^{2}$ 1 1 0.99 1 1 0.99 0.99 0.99
MMRE 0.17 0.11 0.14 0.09 0.09 0.09 0.09 0.06
RMSE 0.02 0.02 0.03 0.02 0.02 0.02 0.03 0.02
Shape* RT RT RT RT RT RT RT RT
LSC $R^{2}$ 1 1 1 1 1 0.99 1 0.99
MMRE 0.18 0.10 0.13 0.11 0.11 0.09 0.08 0.09
RMSE 0.02 0.02 0.02 0.03 0.02 0.02 0.02 0.02
Shape* Q RT RT E Q RT RT E
IDC $R^{2}$ 1 1 1 1 1 1 1 1
MMRE 0.02 0.08 0.11 0.21 0.14 0.13 0.10 0.04
RMSE 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.01
Shape* L RT RT L L Q Q RT
*Selected regression model shapes include Linear (L), Exponential (E), Quadratic (Q),
and Regression Tree (RT)
@@
(g) $𝑆𝑢𝑏𝑗𝑒𝑐𝑡𝑆7{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{Subject\ S7}}$
(h) $𝑆𝑢𝑏𝑗𝑒𝑐𝑡𝑆8{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{Subject\ S8}}$
@@
Figure 4: Selected regression models across subjects using MS based on the training set
V-A2 RQ1.2
Based on which adequacy metrics can an accurate regression model be built to predict FDR from AS based on the training set?
As described in the TEASMA’s procedure in Section III, for each subject, multiple regression models are examined starting with a simple linear regression model and then using more complex non-linear models (quadratic, exponential, and regression trees) when needed. As outlined in Algorithm 1, TEASMA selects by default the regression model with the highest $R^{2}$ as the best model. In Table III, we report detailed results for the selected regression model including $R^{2}$, average RMSE, and MMRE across all $5$-folds, along with the model shape for all subjects. Further, to provide a full report, the details of all examined regression models across all subjects are reported in our replication package [44].
The results for MS show a minimum $R^{2}$ value of 0.98, while the other three metrics show a minimum $R^{2}$ value of 0.99 across all subjects, indicating that all these regression models explain most of the variance in FDR. The average RMSE, across all $K$-folds for all subjects, ranges from 0.01 to 0.05 for MS, from 0.02 to 0.03 for DSC and LSC, and from 0.01 to 0.02 for IDC, indicating high prediction accuracy. Additionally, the average MMRE ranges from 0.04 to 0.25 for MS, from 0.06 to 0.17 for DSC, from 0.08 to 0.18 for LSC, and from 0.02 to 0.21 for IDC, indicating a low relative error. These results suggest that all four investigated metrics can be used to very accurately predict FDR in the training set through (non-linear) regression analysis.
Table III also reports the shape of the regression model using all four adequacy metrics across all subjects. In the case of MS, only S7 exhibits linearity, the shape of S4 is exponential, and regression trees are the best for all other subjects. However, when using SC metrics, the selected regression model consistently has a non-linear shape, with regression trees emerging as the best for all subjects when using DSC. For LSC, the regression models have an exponential shape for two subjects, quadratic for another two, and for the remaining subjects, regression trees are the best. Regarding IDC, four subjects exhibit a linear shape while regression trees are the best for the remaining subjects. As we can see, there is a wide variation of shapes across metrics and subjects.
@@
This suggests that across all metrics, standard regression functions are often not a good fit and regression trees offer more flexibility by optimally partitioning the AS range and computing averages within partitions. To prevent overfitting of the regression trees, which would yield poor results on test sets, we limited their maximum depth to 5.
To visualize our results, we present scatterplots and fitted regression curves between FDR and AS with 95% prediction intervals (PI) for each subject, based on all input subsets sampled from the training set. We compute PIs in linear, quadratic, and exponential regressions using a standard method based on the fitted regression models, and for the regression tree models, we compute PIs based on a bootstrapping method for non-parametric regression models [48]. PIs depict, for each specific AS value, the likely range of potential FDR values for a 95% confidence level. The regression models built using all metrics display narrow PIs across all subjects, suggesting that these models can be confidently used to predict the FDR of a test set with limited uncertainty.
@@
The fitted regression curves for MS across all subjects are displayed in Figure 4. For all subjects, except S7, where the shape is highly non-linear, as subset size increases, MS initially shows rapid growth compared to FDR, indicating that the first mutants are relatively easy to kill when compared to actual faults. But then the rate of increase in MS starts to quickly slow down over time, suggesting that mutants are rapidly becoming more difficult to kill. As illustrated in Figure 4, identical MS values can result in very different levels of fault detection abilities across subjects. For instance, in subject S6, an MS of 0.8 is associated with an FDR of 0.2 while in subject S4, the same MS is associated with an FDR of 0.8. The only exception is subject S7, where the relationship between FDR and MS is linear. However, the slope of the regression line is slightly greater than one, suggesting that killing mutants is harder than detecting faults. These observations confirm that, though mutants may not be representative of faults, it does not prevent us from accurately modeling their relationship.
Due to space constraints, the fitted regression curves for other metrics are provided in our replication package [44]. However, we report and discuss our observations here. When using both SC metrics, DSC, and LSC, the shape of the selected regression model is non-linear across all subjects. However, when using DSC, while the maximum FDR is consistently achieved across all subjects, the maximum DSC coverage (close to 1) is only achieved for subject S8. Moreover, the coverage of subsets with FDR close to 1 varies across subjects and is less than 0.4 for S6 and S7, less than 0.6 for S1, S2, and S5, and less than 0.7 for S3 and S4. For LSC, the maximum coverage is achieved on three subjects (S2, S4, and S6) but the coverage of subsets with FDR close to 1 varies across subjects. This indicates that achieving maximum coverage based on DSC and LSC is harder than detecting all faults. As a result, these metrics cannot provide a stopping criterion for selecting an adequate test set other than achieving the maximum coverage, which often requires selecting a large number of inputs. Consequently, this further justifies the importance of a test assessment methodology that can accurately predict a test set’s FDR. In general, the highly non-linear relationship between AS and FDR when using MS, DSC, and LSC, on most subjects, indicates that none of these metrics can be directly used to assess test sets and decide on their adequacy.
@@
The shape of the selected regression model between IDC and FDR using the training subsets is linear for three subjects, with a slope that varies across these subjects and is consistently higher than 1. For the remaining subjects, the selected regression models are non-linear. We can therefore draw similar conclusions as above for other metrics. To summarize, the shape of the relationship between AS and FDR, using the investigated metrics, greatly varies across subjects, which is likely due to the characteristics of the DNN models and associated training sets. For example, differences in internal architectures across DNN models determine the distribution of mutants generated by post-training MOs and contribute to such variation in relationships. Though this is not an issue in practice, this highlights the need to determine the best regression model empirically for each DNN model based on its training set, as specified by TEASMA (Section III).
@@
Answer to RQ1.2: An accurate regression model with a minimum $R^{2}=0.98$ (based on cross-validation) can be fitted for each subject based on its training set to accurately capture the relationship between FDR and AS, whether calculated using MS, DSC, LSC, or IDC. The shape of the regression model is often non-linear and greatly varies across subjects.
@@
TABLE IV: Linear correlation between $\widehat{FDR}$ and $ActualFDR$, for all random test subsets, obtained using each adequacy metric across all subjects
Adequacy Accuracy Subjects
metric metric S1 S2 S3 S4 S5 S6 S7 S8
MS (E1) $R^{2}$ 0.97 0.98 0.94 0.97 0.98 0.98 0.99 0.99
Slope 1.21 1.19 1.16 1.23 0.95 1.08 1.18 1.21
DSC $R^{2}$ 0.99 0.95 0.90 0.98 0.90 0.95 0.97 0.98
Slope 1.46 1.61 1.67 0.93 0.81 1.43 0.93 1.23
LSC $R^{2}$ 0.79 0.97 0.90 0.89 0.99 0.94 0.93 0.99
Slope 3.56 2.01 1.64 8.09 1.90 8.32 1.48 1.48
IDC $R^{2}$ 0.99 0.94 0.90 0.98 0.99 0.90 0.97 0.99
Slope 1.45 1.53 1.62 1.48 1.10 1.52 1.45 1.41
@@
(g) $𝑆𝑢𝑏𝑗𝑒𝑐𝑡𝑆7{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{Subject\ S7}}$
(h) $𝑆𝑢𝑏𝑗𝑒𝑐𝑡𝑆8{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{Subject\ S8}}$